OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (5,1) and (-2,3). how do I figure this out? do I have to find the slope first?

OpenStudy (anonymous):

the equation of a line is y=mx+b so find m (slope) using the points given m= y2-y1/x2-x1=(3-1)/(-2-5)=2/-7=-2/7 so you have y=-2/7x+b to find b, plug in either coordinates given for x and y and isolate for b 1=-2/7(5)+b b=1+(10/7) b=17/7 so the equation is y=-2/7x + 17/7

OpenStudy (anonymous):

the slope is -2/7 I think but what do I do after that?

OpenStudy (anonymous):

yes that is the right slope. see my solution above for the rest of the steps ^^^

OpenStudy (anonymous):

forgot the last step. to make it standard form, just make it equal to zero. so y + 2/7x - 17/7 = 0

OpenStudy (anonymous):

I understand it now. Except the answer isn't an option on my homework ?? Is there another way to do this?

OpenStudy (anonymous):

oh all they did on your answer sheet is make everything into whole numbers

OpenStudy (anonymous):

so just multiply your entire equation by 7 to get rid of the fractions: (7)y + (7)2/7x - (7)17/7 = (7)0 (multiply each term by 7) 7y + 2x -17 = 0 (rearrange if needed) 2x + 7y = 17 (same as B)

OpenStudy (anonymous):

Where does the 7 come from?